# Rajiv Raman

### v2

@@ 5.2 Construction of Outerplanar Supports
$C(\mathcal{H})$ is clearly an outerplanar graph. Consider a subgraph $K\in\mathcal{K}$. Since $K$ induces a single run on $C$ and the subgraphs in $\mathcal{H}$ are strict-containment free, it follows that the subgraphs in $\mathcal{H}_{K}$ appear consecutively in lex. cyclic order on $C(\mathcal{H})$ and thus induce a connected subgraph of $C(\mathcal{H})$.
Now assume that subgraphs in $\mathcal{H}$ induce single runs on $C$, and those in $\mathcal{K}$ can have multiple runs. We claim that there is an outerplanar support with outer cycle $C(\mathcal{H})$. But, before we do that, we start with the following simple consequence of $(C,\mathcal{H},\mathcal{K})$ being $axax$-free.
@@
The proof for the moreover part is similar. Indeed, let $h^{\prime}\in\mathrm{arc}[v,u]$ s.t. $h^{\prime}\in H^{\prime}\setminus H$. Also, note that $x\in H^{\prime}\setminus H$. Hence, the subgraphs $H,H^{\prime}$ form an $axax$-pattern as witnessed by the cyclic sequence $h^{\prime},u,x,v$; a contradiction.
@@
Finally, suppose that $K$ intersects both $C_{L}$ and $C_{R}$. Since $(C,\mathcal{K})$ is $axax$-free and $K_{0}$ contains $u_{0}$ and $v_{0}$, it follows by Proposition 19, that $K$ contains either $u_{0}$ or $v_{0}$. Therefore, either $H_{1}$ or $H^{\prime}_{s}$ is in $\mathcal{H}_{K}$. Now, the fact that $\mathcal{H}_{K}$ induces a connected subgraph in $Q$ is identical to the case when $K$ has runs only in $C_{L}$ s.t. $\mathcal{H}_{K}\cap\mathcal{H}_{u_{0}}\neq\emptyset$. Since $K$ was chosen arbitrarily, $Q$ is a support.
@@
Now we prove the result for the general setting i.e., when subgraphs in $\mathcal{H}\cup\mathcal{K}$ induce more than one run on $C$.
See 14
@@
Now suppose that all runs of $K$ lie entirely in $\mathrm{arc}[v_{0},u_{0}]$. If each $H\in\mathcal{H}_{K}$ has its representative in $\mathcal{H}_{L}$, then the fact that $\mathcal{H}_{K}$ induces a connected subgraph of $Q$ follows from the fact that $Q_{L}$ is an intersection support for $(C,\mathcal{H}_{L},\mathcal{K})$. Otherwise, if an $H\in\mathcal{H}_{K}$ does not have its representative in $\mathcal{H}_{L}$, then by construction of $\mathcal{H}_{L}$, $H\cap\mathrm{arc}(u_{0},v_{0})\neq\emptyset$. By Proposition 19 therefore, $H$ contains $u_{0}$ or $v_{0}$ and each run of $H$ in $\mathrm{arc}[v_{0},u_{0}]$ is contained in a run of $H_{0}$. It follows that $K\cap H_{0}\neq\emptyset$ and thus $K\cap H^{\prime\prime}_{0}\neq\emptyset$. Recall that the subgraph in $\mathcal{H}$ containing $u_{0}$ or those containing $v_{0}$ appear consecutively in the outer cycle $C(\mathcal{H}_{R})$ of $Q_{R}$, and $K$ intersects a prefix of these sequences of subgraphs. Since $K$ intersects $H_{0}$ which contains both $u_{0}$ and $v_{0}$, $\mathcal{H}_{K}$ induces a connected subgraph of $Q$.
Finally, let $K$ intersect both $\mathrm{arc}[u_{0},v_{0}]$ and $\mathrm{arc}[v_{0},u_{0}]$. Note that in this case $K$ intersects $H_{0}$ (and hence with $H^{\prime}_{0}$ and $H^{\prime\prime}_{0}$) since $u_{0},v_{0}\in H_{0}$ and $(C,\mathcal{H},\mathcal{K})$ is strong $axax$-free. Each $H\in\mathcal{H}\setminus\{H_{0}\}$ has a representative vertex in one of $Q_{L}$ or $Q_{R}$ which are supports for the derived cycle systems with $H_{0}$ in both $Q_{L}$ and $Q_{R}$. It follows that $\mathcal{H}_{K}$ induces a connected subgraph of $Q$. Since $K$ was chosen arbitrarily, $Q$ is a support.
@@ 5.3 Implementation
In this section, we show a polynomial running time of our algorithm to construct an outerplanar support. First, we give a short argument that $|\mathcal{H}|$ and $|\mathcal{K}|$ are polynomially bounded. Possibly the bound here is not sharp, but is sufficient for our purposes since we only want to establish a polynomial running time. We use the results of [ackerman2020coloring] and [RR18].
@@
Hence, the VC-dimension of the set system $(C,\mathcal{H})$ is at most 4, and this implies, by the Sauer-Shelah lemma (See Chapter 10 in [Mat02LecDiscGeom]) that $|\mathcal{H}|,|\mathcal{K}|=O(n^{4})$.
@@
Let $(C,\mathcal{H},\mathcal{K})$ be the resulting strong $axax$-free cycle system. Then $|C|=O(n)$, and by Theorem 22, $|\mathcal{H}|,|\mathcal{K}|=O(n^{4})$. We only need to show that our algorithm in Theorem 14 runs in time $O(n^{6})$.
@@
In the general case, a subgraph in $\mathcal{H}$ can have multiple runs. To compute an $H\in\mathcal{H}$ minimizing $\ell(H)$, we can store the subgraphs in $\mathcal{H}$ in a heap ordered by $\ell(H)$. We can split the problem into two sub-problems in $O(|C|\max\{|\mathcal{H}|,|\mathcal{K}|\})$ time. Since we add at most $|C|-3$ chords, the overall running time is bounded above by $O(|C|^{2}\max\{|\mathcal{H}|,|\mathcal{K}|\})$. Hence, the overall running time of the algorithm is $O(n^{6})$.
Outerplanar graphs have treewidth at most $2$ and the above result shows that if the subgraphs are non-piercing, then the VC dimension of the set system defined by $(C,\mathcal{H})$ is at most $4$. We extend this result in Section 6 to graphs of bounded treewidth and we show that for a non-piercing graph system $(G,\mathcal{H})$ of treewidth $t$, the VC-dimension of set system defined by $(V(G),\mathcal{H})$ is at most $33t+$ 3.
@@ 5.4 Primal and Dual Supports: Revisited
The existence of an intersection support also implies a primal and a dual support. However, we emphasize here that for a primal and a dual support, we require strictly weaker conditions. In particular, for a primal support, we require the cycle system $(C,\mathcal{H})$ to be $abab$-free and not necessarily $axax$-free, and for a dual support, we require $(C,\mathcal{H})$ to be $axax$-free. The result for a primal support directly follows from the following result of Raman and Singh [raman2025supportsgraphsboundedgenus].
@@
For each $v\in V(C)$ such that $c(v)=\mathbf{r}$, we delete it and make its neighbours adjacent. Let $C^{\prime}$ be the resulting cycle obtained form $C$, and =$\mathcal{H}^{\prime}=\{H\cap C^{\prime}:H\in\mathcal{H}\}$. Then $(C^{\prime},\mathcal{H}^{\prime})$ is an $abab$-free cycle system. By Lemma 24, there is a set $D$ of non-intersecting chord added to $C^{\prime}$ such that each subgraph in $\mathcal{H}^{\prime}$ induces a connected subgraph of $C^{\prime}\cup D$. Thus $C^{\prime}\cup D$ is an outerplanar primal support for $(C^{\prime},\mathcal{H}^{\prime})$. Note that this is also a primal support for $(C,\mathcal{H})$.
@@
The reason why $(G,\mathcal{H})$ does not admit a dual outerplanar support is that the induced system $(C,\mathcal{H})$ is not $axax$-free - $H_{4}$ forms an $axax$-pair with each of $H_{1},H_{2}$ and $H_{3}$. However, if $axax$-freeness is a sufficient condition for $(C,\mathcal{H})$ to exhibit an outerplanar dual support and it follows from Theorem 14.
@@
Let $\mathcal{K}=\{\{v\}:v\in V(C)\}$ be a collection of subgraphs each consisting of a single vertex. Then $(C,\mathcal{K})$ is $axax$-free. Therefore, the intersection system $(C,\mathcal{H},\mathcal{K})$ is $axax$-free. Note that the strong $axax$-free property is trivially satisfied by $(C,\mathcal{H},\mathcal{K})$. By Theorem 14, it has an intersection support for $Q$, which is also a dual support for $(C,\mathcal{H})$ since each $v\in V(C)$ corresponds to a subgraph in $K\in\mathcal{K}$.
@@
Figure 7: An example of an $axax$-free, but not strong $axax$-free system $(C,\mathcal{H,K})$ that does not admit an outerplanar support.
6 Support of Bounded Treewidth
Lemma 27.
Proof.
The proofs follow immediately from the fact that $\mathcal{H}$ are non-piercing and the adhesion set corresponding to any edge in $E(T)$ is a separator in $G$.
Lemma 28.
Proof.
Since $G$ is chordal, each bag induces a complete subgraph of $G$. Hence, each $H_{x}\in\mathcal{H}_{x}$ is connected. If $x$ is not a root of $T$, let $y$ be the parent of $x$. Then, for $H_{x},H^{\prime}_{x}\in\mathcal{H}_{x}$, the subgraph $H_{x}\setminus H^{\prime}_{x}$ of $G_{x}$ is connected since $H\setminus H^{\prime}$ is connected and $A_{xy}$ is a separator in $G$. Therefore, $(G_{x},\mathcal{H}_{x})$ is a non-piercing graph system. Finally, that $(T_{x},\mathcal{B}_{x})$ is a tree decomposition, is straightforward.
6.1 Primal Support
Lemma 29.
Proof.
To show that $Q$ is a support, consider an $H\in\mathcal{H}$. Since $H$ is a connected subgraph in $G$, the vertices in $V(H)$ lie in bags corresponding to a connected sub-tree of $T$. Further, as $(T,\mathcal{B})$ is an easy tree decomposition, $\mathbf{b}(V(H))$ lie in bags of $\mathcal{B}^{\prime}$ corresponding to a connected sub-tree of $T$. Since each $B^{\prime}\in\mathcal{B}^{\prime}$ induces a complete subgraph of $Q$, it implies $V(H)$ induces a connected subgraph of $Q$.
Lemma 30.
Let $(G,\mathcal{H})$ be a non-piercing graph system with $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$. A tree decomposition $(T,\mathcal{B})$ of $G$ of width $t$ can be transformed into a tree decomposition $(T,\mathcal{B}^{\prime})$ of width at most $2^{t+2}+t$.
Proof.
We claim that $(T,\mathcal{B}^{\prime})$ is an easy tree decomposition. Suppose not. Let $H\in\mathcal{H}$ s.t. $H\cap B_{\rho}\neq\emptyset$ and $H\cap\mathbf{b}(B^{\prime}_{\rho})=\emptyset$. Then, $H\cap A_{x\rho}\neq\emptyset\neq H\cap A_{y\rho}$ by assumption. Let $H\cap A_{x\rho}=S$ and $H\cap A_{y\rho}=S^{\prime}$. Let $H^{\prime}$ be the minimal subgraph in $\mathcal{M}_{S}$ whose blue vertex was added to $B^{\prime}_{\rho}$ and let $H^{\prime\prime}$ be the minimal subgraph in $\mathcal{M}_{S^{\prime}}$ whose blue vertex was added to $B^{\prime}_{\rho}$. Then, $H$ and $H^{\prime}$ are incomparable in $(G_{x},\preceq)$. By part $(ii)$ of Lemma 27, this implies $H$ and $H^{\prime}$ are identical in $G_{-x}$. Similarly, $H$ and $H^{\prime\prime}$ are incomparable in $(G_{y},\preceq)$ and by Lemma 27 $H$ and $H^{\prime\prime}$ are identical in $G_{-y}$. It follows that $H^{\prime}|_{x}$ and $H^{\prime\prime}|_{x}$ are incomparable in $G_{x}$, and $H^{\prime}|_{y}$ and $H^{\prime\prime}|_{y}$ are incomparable in $G_{y}$. Also, note that $H^{\prime\prime}\cap A_{x\rho}=S=H^{\prime}\cap A_{x\rho}$. But this implies $H^{\prime}\setminus H^{\prime\prime}$ has at least two connected components - one in $G_{x}$ and one in $G_{y}$ contradicting the assumption that the subgraphs in $\mathcal{H}$ are non-piercing. Hence, $(T,\mathcal{B}^{\prime})$ is an easy tree decomposition.
See 7
Proof.
Let $(T,\mathcal{B})$ be a tree decomposition of $G$ of width $t$ where $T$ is a binary rooted tree. If $(T,\mathcal{B})$ is an easy tree decomposition, we are done by Lemma 29. Otherwise, we transform it into an easy tree decomposition of width at most $2^{t+2}+t$ by Lemma 30. We then obtain a primal support $Q$ of the same treewidth by Lemma 29.
6.2 Dual Support
Lemma 31.
Proof.
For a vertex $v\in V$, consider $\mathcal{H}_{v}$, the set of all subgraphs containing $v$. Since $(T,\mathcal{B})$ is a tree decomposition, at each node $x\in V(T)$ s.t. $v\in B_{x}$, $\mathcal{H}_{v}\subseteq B^{\prime}_{x}$, where $B^{\prime}_{x}$ is the bag at node $x$ in $\mathcal{B}^{\prime}$, i.e., corresponding to $B_{x}$. Since the vertices in $B^{\prime}_{x}$ are pairwise adjacent, it implies $\mathcal{H}_{v}$ induces a connected subgraph of $Q^{*}$. Hence, $Q^{*}$ is a dual support.
Lemma 32.
Proof.
$|\textrm{unique}(\mathcal{H}^{\prime}_{A_{xz}}|_{x})|<2^{t+1}$ (1)
We now argue that $Q^{*}$ is a dual support for $(G,\mathcal{H})$. Consider a vertex $v\in V(G)$. The algorithm above ensures that each subgraph is pushed at most once. Suppose $H^{\prime}\in\mathcal{H}_{v}$ was pushed by $H$ so that its modified copy $H^{\prime\prime}=H^{\prime}\setminus H$ does not cover $v$. Then, $H\in\mathcal{H}_{v}$. Let $e=\{a,b\}$ be the connecting edge between $H$ and $H^{\prime\prime}$ such that $a\in H$ and $b\in H^{\prime\prime}$. Since $(T,\mathcal{B})$ is a valid tree-decomposition, there is a bag $B$ containing both $a$ and $b$. Let $H_{1}$ be the unique representative of $H^{\prime\prime}$ in $\textrm{unique}(\mathcal{H}^{\prime})$. Since Algorithm $k$-SDS puts a complete graph on the subgraphs intersecting $B$, it implies $H$ and $H_{1}$ are adjacent in $Q^{\prime}$. In $Q^{*}$, we made $H^{\prime}$ adjacent to $H_{1}$. Hence, $Q^{*}$ is a dual support for $(G,\mathcal{H})$.
See 8
Proof.
Let $(T,\mathcal{B})$ be a tree decomposition of $G$ of width $t$. If $(T,\mathcal{B})$ is $2^{4(t+1)}$-sparse, then we are done. Otherwise, by Lemma 32, we obtain a dual support $Q^{*}$ for $(G,\mathcal{H})$, of treewidth at most $2^{4(t+1)}$.
6.3 Intersection Support
Lemma 33.
Proof.
See 9
Proof.
Since the subgraphs in $\mathcal{H}$ are non-piercing, by Lemma 32, we can obtain a collection $\mathcal{H}^{\prime}$ s.t. $(T,\mathcal{B}^{\prime})$ is $2^{4(t^{\prime}+1)}$-sparse with respect to $\textrm{unique}(\mathcal{H}^{\prime})$. Note that $(T,\mathcal{B}^{\prime})$ remains $\mathcal{K}$-easy w.r.t. the intersection system $(G,\textrm{unique}(\mathcal{H}^{\prime}),\mathcal{K})$. Now, by Algorithm $k$-SDS we obtain a tree decomposition of a graph $Q^{*}$ that by Lemma 31 is a dual support for $(G,\textrm{unique}(\mathcal{H}^{\prime}))$ and $\textsc{tw}(Q^{*})\leq 2^{4(t^{\prime}+1)}=2^{{2^{t+4}}+4(t+1)}$. Since $Q^{*}$ is obtained by $k$-SDS, it also an intersection support for $(G,\textrm{unique}(\mathcal{H}^{\prime}),\mathcal{K})$. By Theorem 8, we can extend $Q^{*}$ to a dual support $\tilde{Q}$ for $(G,\mathcal{H})$ such that $\textsc{tw}(\tilde{Q})=\textsc{tw}(Q^{*})$. Then $\tilde{Q}$ is also an intersection support for $(G,\mathcal{H},\mathcal{K})$ since the tree decomposition $(T,\mathcal{B}^{\prime})$ is $\mathcal{K}$-easy.
6.4 Implementation
In this section, we show that our algorithms for the construction of supports of bounded treewidth run in polynomial time if the treewidth of the host graph is bounded. We argue by showing a result below that generalizes Theorem 22 to graphs of treewidth $t$. We show that if the treewidth of $G$ is bounded, then $|\mathcal{H}|$ and $|\mathcal{K}|$ are bounded by $\textrm{poly}(n)$ where $n=|V(G)|$.
Theorem 34.
Proof.
Next, we define a set of subgraphs in $\mathcal{H}$ that intersect with both $S_{1}$ and $S_{2}$. Such a set of subgraphs must exist since $S$ is shattered by element of $\mathcal{H}$. Let $J_{i}\in\mathcal{H}$ be a subgraph such that $J_{i}\cap S=H_{i}\cup H^{\prime}_{i}$ for $i=1,2,\ldots k$. Then, each $J_{i}$ contains exactly $2\lfloor\frac{t+2}{2}\rfloor$ vertices of $S$ and hence, they are pairwise incomparable in $S$. Also, each $J_{i}$ intersects the adhesion set $A_{e}$ as they are connected subgraphs. Since the subgraphs in $\mathcal{H}$ are non-piercing, by part $(ii)$ of Lemma 27, $J_{i}$ and $J_{\ell}$ should be incomparable in $A_{e}$ for $i\neq\ell$. But there are only $\binom{t+1}{\lfloor\frac{t+1}{2}\rfloor}<k$ subsets of $A_{e}$ that are pairwise incomparable. This contradicts the fact that the set $S$ can be shattered. Hence, the VC-dimension of $(V(G),\mathcal{H})$ is at most $33t+$ 3. By Sauer-Shelah lemma (see [Mat02LecDiscGeom], Chap. 10) therefore, $|\mathcal{H}|\leq O(n^{3t+3})$.
Now, we are ready to show the running time of our algorithms.
Theorem 35.
Proof.
7 Lower Bounds
In this section, we show that there exist graphs of treewidth $t$ whose (primal or dual) support requires treewidth $\Omega(2^{t})$.
Theorem 36.
There is a non-piercing graph system $(G,\mathcal{H})$ with $c:V(G)\to\{\mathbf{r},\mathbf{b}\}$ s.t. $\textsc{tw}(G)\leq n+1$, but $\textsc{tw}(Q)\geq 2^{\lceil n/2\rceil}$ for any primal support $Q$.
Proof.
Let $m=\lceil n/2\rceil$ and $N=2^{m}$. We construct a graph $G$ and non-piercing subgraphs $\mathcal{H}$ as follows. Let $B=\{b_{i,j}:i,j\in\{1,\ldots,N\}\}$ be an $N\times N$ grid of isolated vertices colored $\mathbf{b}$ such that there is no edge between the vertices of $B$. Let $C=\{c_{1},\ldots,c_{m}\}$ and $R=\{r_{1},\ldots,r_{m}\}$ be two sets of vertices colored $\mathbf{r}$. Let $\textrm{bin}(x)$ denote the binary representation of $x\in\mathbb{N}$. By $\textrm{bin}(x)$ of an indexed set $X$, we mean the subset of $X$ with indices corresponding to $1$s in $\textrm{bin}(x)$. Since $\log N=m$, for any integer $0\leq x\leq N$, $\textrm{bin}(x)$ of $C$ and $\textrm{bin}(x)$ of $R$ are well-defined.
For a horizontal pair of grid vertices $e=\{b_{i,j},b_{i,j+1}\}$, we connect these vertices to $\textrm{bin}(j)$ of $C$, $\textrm{bin}(j+1)$ of $C$ and $\textrm{bin}(i)$ of $R$. This gives an induced subgraph $H_{i,j,i,j+1}$. Similarly, for a vertical pair $e^{\prime}=\{b_{i^{\prime},j^{\prime}},b_{i^{\prime}+1,j^{\prime}}\}$, we connect these vertices to $\textrm{bin}(i^{\prime})$ of $R$, $\textrm{bin}(i^{\prime}+1)$ of $R$, and $\textrm{bin}(j^{\prime})$ of $C$. This gives us an induced subgraph $H_{i^{\prime},j^{\prime},i^{\prime}+1,j^{\prime}}$. This completes simultaneously, the construction of $G$ and the subgraphs $\mathcal{H}$.
We claim that $(G,\mathcal{H})$ is a non-piercing graph system of treewidth at most $n+1$. The graph $G$ thus constructed is a bipartite graph with bipartition $(B,R\cup C)$, since the vertices in $B$ are pairwise non-adjacent, and the vertices in $R\cup C$ are pairwise non-adjacent. Since $|R\cup C|=2\left\lceil n/2\right\rceil$, it implies that $\textsc{tw}(G)\leq n+1$.
The subgraphs in $\mathcal{H}$ are connected. Let $H,H^{\prime}\in\mathcal{H}$ be arbitrary. Since they correspond to distinct pairs of points of the grid, they differ in at least one grid vertex, and they also differ in their adjacency in either $R$ or in $C$. Hence, $H\setminus H^{\prime}$ is connected via the grid vertex in $H\setminus H^{\prime}$. Similarly, $H^{\prime}\setminus H$ is connected. Therefore, $\mathcal{H}$ is a collection of non-piercing subgraphs of $G$.
Each $H\in\mathcal{H}$ contains exactly two vertices of $B$, which should be adjacent in any primal support for $(G,\mathcal{H})$. Therefore, any primal support $Q$ must contain a grid graph on $B$. Hence, $\textsc{tw}(Q)\geq N$.
Theorem 37.
There is a non-piercing graph system $(G,\mathcal{H})$ s.t. $\textsc{tw}(G)\leq n+1$ and for any dual support $Q^{*}$, $\textsc{tw}(Q^{*})\geq 2^{\lceil n/2\rceil}-1$.
Proof.
The construction is similar to the construction of the primal system. Let $m=\lceil n/2\rceil$ and $N=2^{m}$. Let $B=\{b_{i,j}:i,j\in\{1,\ldots,N\}\}$ be an $N\times N$ grid of isolated vertices. Let $C=\{c_{1},\ldots,c_{m}\}$ and $R=\{r_{1},\ldots,r_{m}\}$ be two sets of vertices, none of which are pairwise adjacent.
The graph system $(G,\mathcal{H})$ is constructed as follows. A face $F_{i,j,i+1,j+1}$ for $i,j\in\{1,\ldots,N-1\}$ is the subset $\{b_{i,j},b_{i,j+1},b_{i+1,j},b_{i+1,j+1}\}$ of $B$. For each $F_{i,j,i+1,j+1}$, we connect all its vertices to $\textrm{bin}(i)$ and $\textrm{bin}(i+1)$ of $R$, and $\textrm{bin}(j)$ and $\textrm{bin}(j+1)$ of $C$. Finally, for every two consecutive vertices on the grid, we put a vertex in the middle. Let $B^{\prime}$ be the set of vertices added. Thus, $B\cup B^{\prime}$ forms a grid of size $(2N-1)\times(2N-1)$. For each vertex of $B^{\prime}$ that lies between two consecutive vertices $x,y$ of $B$ row-wise or column-wise, we make it adjacent to all the vertices in $R\cup C$ that are adjacent to both $x$ and $y$. This completes the construction of $G$. Now, for each face $F_{i,j,i+1,j+1}$, we add an induced subgraph $H_{i,j,i+1,j+1}$ to $\mathcal{H}$ consisting of the vertices of $F_{i,j,i+1,j+1}$, the four vertices of $B^{\prime}$ that lie between the vertices of this face, and the common neighbours of these 8 vertices in $R\cup C$ i.e., $\textrm{bin}(i)$ and $\textrm{bin}(i+1)$ of $R$, and $\textrm{bin}(j)$ and $\textrm{bin}(j+1)$ of $C$. This results in connected induced subgraphs $\mathcal{H}$ of $G$.
$G$ is a bipartite graph with bipartition $(B,C\cup R)$. Since $|C\cup R|=2m$, $\textsc{tw}(G)\leq 2m$. Each subgraph in $\mathcal{H}$ contains 8 vertices of $B\cup B^{\prime}$. Since any two subgraphs $H,H^{\prime}\in\mathcal{H}$ share at most 3 vertex of $B\cup B^{\prime}$ along a row or column of the grid. Hence, $H\setminus H^{\prime}$ is connected via the vertices of $B\cup B^{\prime}$ in $H\setminus H^{\prime}$. Thus, $\mathcal{H}$ is a collection of non-piercing subgraphs.
Since each vertex in $B^{\prime}$ lies in exactly two subgraphs, the subgraphs in $\mathcal{H}$ form a grid of size $(N-1)\times(N-1)$. Therefore, for any dual support $Q^{*}$, $\textsc{tw}(Q^{*})\geq N-1$.

### v3

@@ 7.2 Construction of Outerplanar Supports
$C(\mathcal{H})$ is clearly an outerplanar graph. Consider a subgraph $K\in\mathcal{K}$. Since $K$ induces a single run on $C$ and the subgraphs in $\mathcal{H}$ are strict-containment free, it follows that the subgraphs in $\mathcal{H}_{K}$ appear consecutively in lex. cyclic order on $C(\mathcal{H})$ and thus induce a connected subgraph of $C(\mathcal{H})$. ∎
Now, assume that subgraphs in $\mathcal{H}$ induce single runs on $C$, and those in $\mathcal{K}$ can have multiple runs. We claim that there is an outerplanar support with outer cycle $C(\mathcal{H})$. But, before we do that, we start with the following simple consequence of $(C,\mathcal{H},\mathcal{K})$ being $axax$-free.
@@
The proof for the moreover part is similar. Indeed, let $h^{\prime}\in\mathrm{arc}[v,u]$ s.t. $h^{\prime}\in H^{\prime}\setminus H$. Also, note that $x\in H^{\prime}\setminus H$. Hence, the subgraphs $H,H^{\prime}$ form an $axax$-pattern as witnessed by the cyclic sequence $h^{\prime},u,x,v$; a contradiction. ∎
@@
Finally, suppose that $K$ intersects both $C_{L}$ and $C_{R}$. Since $(C,\mathcal{K})$ is $axax$-free and $K_{0}$ contains $u_{0}$ and $v_{0}$, it follows by Proposition 31, that $K$ contains either $u_{0}$ or $v_{0}$. Therefore, either $H_{1}$ or $H^{\prime}_{s}$ is in $\mathcal{H}_{K}$. Now, the fact that $\mathcal{H}_{K}$ induces a connected subgraph in $Q$ is identical to the case when $K$ has runs only in $C_{L}$ s.t. $\mathcal{H}_{K}\cap\mathcal{H}_{u_{0}}\neq\emptyset$. Since $K$ was chosen arbitrarily, $Q$ is a support. ∎
@@
Now, we prove the result for the general setting i.e., when subgraphs in $\mathcal{H}\cup\mathcal{K}$ induce more than one run on $C$.
See 26
@@
Now, suppose that all runs of $K$ lie entirely in $\mathrm{arc}[v_{0},u_{0}]$. If each $H\in\mathcal{H}_{K}$ has its representative in $\mathcal{H}_{L}$, then the fact that $\mathcal{H}_{K}$ induces a connected subgraph of $Q$ follows from the fact that $Q_{L}$ is an intersection support for $(C,\mathcal{H}_{L},\mathcal{K})$. Otherwise, if an $H\in\mathcal{H}_{K}$ does not have its representative in $\mathcal{H}_{L}$, then by construction of $\mathcal{H}_{L}$, $H\cap\mathrm{arc}(u_{0},v_{0})\neq\emptyset$. By Proposition 31 therefore, $H$ contains $u_{0}$ or $v_{0}$ and each run of $H$ in $\mathrm{arc}[v_{0},u_{0}]$ is contained in a run of $H_{0}$. It follows that $K\cap H_{0}\neq\emptyset$ and thus $K\cap H^{\prime\prime}_{0}\neq\emptyset$. Recall that the subgraph in $\mathcal{H}$ containing $u_{0}$ or those containing $v_{0}$ appear consecutively in the outer cycle $C(\mathcal{H}_{R})$ of $Q_{R}$, and $K$ intersects a prefix of these sequences of subgraphs. Since $K$ intersects $H_{0}$ which contains both $u_{0}$ and $v_{0}$, $\mathcal{H}_{K}$ induces a connected subgraph of $Q$.
Finally, let $K$ intersect both $\mathrm{arc}[u_{0},v_{0}]$ and $\mathrm{arc}[v_{0},u_{0}]$. Note that in this case $K$ intersects $H_{0}$ (and hence with $H^{\prime}_{0}$ and $H^{\prime\prime}_{0}$) since $u_{0},v_{0}\in H_{0}$ and $(C,\mathcal{H},\mathcal{K})$ is strong $axax$-free. Each $H\in\mathcal{H}\setminus\{H_{0}\}$ has a representative vertex in one of $Q_{L}$ or $Q_{R}$ which are supports for the derived cycle systems with $H_{0}$ in both $Q_{L}$ and $Q_{R}$. It follows that $\mathcal{H}_{K}$ induces a connected subgraph of $Q$. Since $K$ was chosen arbitrarily, $Q$ is a support. ∎
@@ 7.3 Implementation
In this section, we show a polynomial running time of our algorithm to construct an outerplanar support. First, we give a short argument that $|\mathcal{H}|$ and $|\mathcal{K}|$ are polynomially bounded. To show this, we again use an argument based on VC-dimension, together with results of [ackerman2020coloring] and [RR18]. Possibly the bound here is not sharp, but it is sufficient for our purposes since we only want to establish a polynomial running time. We start with the definition of an $abab$-free cycle system.
Definition 34 ($abab$-free).
Note that, if $(C,\mathcal{H})$ is $axax$-free, then it is $abab$-free. The notion of cycle $abab$-free is equivalent to that of $ABAB$-free hypergraphs introduced by Ackerman et al. [ackerman2020coloring]. For the ease of exposition, we use $abab$-free instead of $ABAB$-free in the proof of Theorem 35 below.
@@
Hence, the VC-dimension of the set system $(C,\mathcal{H})$ is at most 4, and this implies, by the Sauer-Shelah lemma (See Chapter 10 in [Mat02LecDiscGeom]) that $|\mathcal{H}|=O(n^{4})$. Similarly, $|\mathcal{K}|=O(n^{4})$. ∎
@@
Let $(C,\mathcal{H},\mathcal{K})$ be the resulting strong $axax$-free cycle system. Then $|C|=n$, and by Theorem 35, $|\mathcal{H}|,|\mathcal{K}|=O(n^{4})$. We only need to show that our algorithm in Theorem 26 runs in time $O(n^{6})$.
@@
In the general case, a subgraph in $\mathcal{H}$ can have multiple runs. To compute an $H\in\mathcal{H}$ minimizing $\ell(H)$, we can store the subgraphs in $\mathcal{H}$ in a heap ordered by $\ell(H)$. We can split the problem into two sub-problems in $O(|C|\max\{|\mathcal{H}|,|\mathcal{K}|\})$ time. Since we add at most $|C|-3$ chords, the overall running time is bounded above by $O(|C|^{2}\max\{|\mathcal{H}|,|\mathcal{K}|\})$. Hence, the overall running time of the algorithm is $O(n^{6})$. ∎
@@ 7.4 Primal and Dual Supports: Revisited
The existence of an intersection support also implies a primal and a dual support. However, we emphasize here that for a primal and a dual support, we require strictly weaker conditions. In particular, for a primal support, we require the cycle system $(C,\mathcal{H})$ to be $abab$-free and not necessarily $axax$-free, and for a dual support, we require $(C,\mathcal{H})$ to be $axax$-free. The result for a primal support directly follows from the following lemma of Raman and Singh [raman2025supportsgraphsboundedgenus].
@@
For each $v\in V(C)$ such that $c(v)=\mathbf{r}$, we delete it and make its neighbours adjacent. Let $C^{\prime}$ be the resulting cycle obtained form $C$, and =$\mathcal{H}^{\prime}=\{H\cap C^{\prime}:H\in\mathcal{H}\}$. Then $(C^{\prime},\mathcal{H}^{\prime})$ is an $abab$-free cycle system. By Lemma 37, there is a set $D$ of non-intersecting chord added to $C^{\prime}$ such that each subgraph in $\mathcal{H}^{\prime}$ induces a connected subgraph of $C^{\prime}\cup D$. Thus $C^{\prime}\cup D$ is an outerplanar primal support for $(C^{\prime},\mathcal{H}^{\prime})$. Note that this is also a primal support for $(C,\mathcal{H})$. ∎
@@
The reason why $(G,\mathcal{H})$ does not admit a dual outerplanar support is that the induced system $(C,\mathcal{H})$ is not $axax$-free - $H_{4}$ forms an $axax$-pair with each of $H_{1},H_{2}$ and $H_{3}$. However, $axax$-freeness is a sufficient condition for $(C,\mathcal{H})$ to exhibit an outerplanar dual support and it follows from Theorem 26.
@@
Let $\mathcal{K}=\{\{v\}:v\in V(C)\}$ be a collection of subgraphs each consisting of a single vertex. Then $(C,\mathcal{K})$ is $axax$-free. Therefore, the intersection system $(C,\mathcal{H},\mathcal{K})$ is $axax$-free. Note that the strong $axax$-free property is trivially satisfied by $(C,\mathcal{H},\mathcal{K})$. By Theorem 26, it has an intersection support for $Q$, which is also a dual support for $(C,\mathcal{H})$ since each $v\in V(C)$ corresponds to a subgraph in $K\in\mathcal{K}$. ∎
@@
Figure 8: An example of an $axax$-free, but not strong $axax$-free cycle system $(C,\mathcal{H,K})$ (shown in left), that does not admit an outerplanar support (shown in right).